OpenStudy (anonymous):

4 log x – 6 log (x + 2) ?

OpenStudy (anonymous):

A: \[\frac{ x }{ x+2 }\] B: log x4(x + 2)^6 C: log x (x + 2)^24 D: none of these

OpenStudy (anonymous):

***A: 24 log\[\frac{ x }{ x+2 }\]

OpenStudy (zehanz):

You can use the following rules for logarithms: \[k \cdot \log_{a}p =\log_{a} p^k\]and\[\log_{a}p-\log_{a}q=\log_{a}\frac{ p }{ q } \]

OpenStudy (zehanz):

E.g.: \[4 \log x = \log x^4\]